Need help with review. Please answer both questions. I really need help.

This Question: 1 pt 7 of 8 (6 complete) A sample mean, sample size, and sample standard deviation are provided below. Use the one-mean t-test to perform the required hypothesis test at the 1% significance level. - :28. s:8, n:32, Ho : μ=30. Ha-μ:30 Data Table Ell Click here to view a partial table of values of ta The test statistic ist- (Round to two decimal places as needed) Identify the critical value(s). Select the correct choice below and fill in the answer box within your choice df t0.10 0.050.025 o.0 0.005df 1372 1.812 2.228 2.764 3.16910 11 1.363 1.796 2.201 2.718 3.106 11 12 1.356 1782 2.179 2.681 3.055 12 13 1.350 1.771 2.160 2.650 3.012 13 141 1.345 1.761 2.145 2.624 2.97기 14 151 1341 1753 2131 2.602 294기15 16 1.337 1.746 2.120 2.583 2.921 16 17 1.333 1740 2.110 2.567 2 18 1.330 1.734 2.101 2.552 2.878 18 19 1.328 1.729 2.093 2.539 2.861 19 20 1.325 1.725 2.086 2.528 2.845 20 21 1.323 1.721 2.080 2.518 2.831 2 22 1.321 1.717 2.074 2.508 2.819 22 23 1.319 1.714 2.069 2.500 2.807 23 241 1.318 1.711 2.064 2.492 2.79기 24 25 1.316 1.708 2.060 2.485 2.787 25 26 1.315 1.706 2.056 2.479 2.779 26 27 1.314 1.703 2.052 2.473 2.771 2 28 1.313 1.701 2.048 2.467 2.763 28 29 1.311 1.699 2.045 2.462 2.756 29 30 1.310 1.697 2.042 2.457 2.750 30 31 1.309 1.696 2.040 2.453 2.744 31 32 1.309 1.694 2.037 2.449 2.738 32 33 1.308 1.692 2.035 2.445 2.733 33 34 1.307 1.691 2.032 2.441 2.728 34 35 1.306 1.690 2.030 2.438 2.724 35 361 1.306 1.688 2.028 2.434 2.719 6 37 1.305 1.687 2.026 2431 2.715 37 38| 1.304 1.686 2.024 2.429 2.71238 39 1.304 1.685 2.023 2.426 2.708 39 (Round to three decimal places as needed) A. The critical value is-4° O B. The critical ist C. The critical values are ± ta/,-± 17 the null hypothesis. The data sufficient evidence to conclude that the mean is less than 30 Print Done
Answer #1

7)
a)
t = (Xbar - mu)/(Sd/sqrt(n))
= (28 - 30)/(8/sqrt(32))
= -1.41421
= -1.41

df =n-1 = 31
at alpha = 0.01
P(T < t-critical) = 0.01
t-critical = -2.453

option A) the critical value -t_alpha = -2.453
we fail to reject the null hypothesis
since -1.41 > -2.453

The data does not provide sufficient evidemce